OpenStudy (anonymous):

Score: Which set of side lengths will form a right triangle? 6, 8, 11 10, 24, 25 7, 24, 26 8, 15, 17

OpenStudy (anonymous):

Hint: Pythagorian triplets

OpenStudy (anonymous):

well the greatest side must be the hypotenuse, which means the root of the squared sum of the other two sides must equal the greatest length.. does that help?

OpenStudy (anonymous):

Use Pythagorean Theorem. If \(c^{2} = a^{2} + b^{2}\), then it's a right triangle. If \(c^{2} > a^{2} + b^{2}\), then it's an obtuse triangle. If \(c^{2} < a^{2} + b^{2}\), then it's an acute triangle. Also, Pythagorean Triplets are basically known triangle sides that create a right triangle.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Square the two smaller numbers of the group. Once you've obtained those squares, add them to one another. Square the larger number (the third number) by itself. If the sum of the two squares equals the square of the larger number, that's the group of three that form a right triangle.

OpenStudy (anonymous):

Example: 3, 4, 5 (3)^2 + (4)^2 = 9 + 16 = 25 (5)^2 = 25 Therefore 3,4,5 will form a right triangle.

OpenStudy (anonymous):

Hope that helped.